OpenStudy (anonymous):

prove the identity by changing the left side of the equation to show equivalency to the right side. cos^4 theta - sin^4 theta = 1 - 2 sin^2 theta

OpenStudy (ranga):

a^2 - b^2 = (a+b)(a-b)

OpenStudy (ranga):

cos^4 theta - sin^4 theta = (cos^2 theta + sin^2 theta)(cos^2 theta - sin^2 theta) = (1) * (1 - sin^2 theta - sin^2 theta) = 1 - 2sin^2 theta

OpenStudy (anonymous):

Awesome! Than you!

OpenStudy (anonymous):

*thank